Physician Visits/year

Total Utility

Marginal Utility

1

20

?    

2

30

?   

3

34

?    

4

?   

2

5

?   

1

At 4 physician visits per year, the consumer's total utility is

a. 40
b. 36
c. 34
d. 2
e. None of the above
Answer #1

ANSWER:

Marginal utility is the change in total utility with respect to the no of visits to the physician each year.

physician visits per year total utility marginal utility
1 20 = 20 / 1 = 20
2 30 = (30 - 20) / (2 - 1) = 10
3 34 = (34 - 30) / (3 - 2) = 4
4 = 34 + 2 = 36 2
5 = 36 + 1 = 37 1

so the correct answer is option b that is 36.